OpenStudy (anonymous):

log7(1 - x) = 3 with 7 being the base of the log

OpenStudy (cwrw238):

(1- x) = 7^3

OpenStudy (ash2326):

We have \[\log_7 (1-x)=3\] Raise both sides by power of 7 we get \[7^{\log_7 (1-x)}=7^3\] we know \[a^{\log_a x}=x\] so \[1-x=7^3\] I think you can solve now!!